MINDTA Chapt Monroe Manufacturing has a quick ratio of 2.00x, $25,425 in cash, $14,125 in accounts receivable, some inventory, total current assets of and total current liabilities of $19,775. The company reported annual sales of $200,000 in the most recent annual report. Over the past year, how often did Monroe Manufecturing sell and replace its inventory? O 286x 8.01x 12.98x O 11.80 x The inventory turnover ratio across companies in the manufacturing industry is 10.03x. Based on this information, which of the following statements is true for Monroe Manufacturing? O Monroe Manufacturing is holdng more inventory per dolar of sales compared to the industry average O Monroe Manufacturing is holding less inventory per dellar of sales compared to the industry average. toys-Like Games Inc. and Our Play Inc. Like Games was launched eight years two years. However, both companies have an You are analyzing two companies that manufacture electronikc ago, whereas Our Play is a relatively new company that has been in equal market share with sales of $200,000 each. Youve collected sales for all industry competitors was $510,000. As an analyst, you want to make comments on the companies in the coming year. Youve collected data from the companies financial statements. This information is listed as follows: company data to compare Like Games and Our Play. Last year, the average want to make comments on the expected performance of these two Data Collected (in dollars) Our Play 7,800 Industry Average 7,700 433,500 469,200 Accounts receivable 5,400 110,000 190,000 Total assets 250,000 Using this information, complete the following statements to include in your analysis. 1. Our Play has tme to collect cash from its customers than it takes Like Games. 2. Like Gamess fixed assets turnover ratio is acquisition coat of ts fixed assets is recorded at historic values when the company bought its assets and has been deprecisted since then Assuming that sixed assets prices (not book values) rose over the past six years due to inflation, Our Play paid a days of sales tied up in receivables, which is much than the industry average. It takes Our Play than that of Our Play. This is because Like Games was fermed eight years ago, so the amount for its fixed O Type here to search PrtScn Hom 4
s holding more inventory per dollar of sales compared to the industry average. ng is hoiding less inventory per dollar of sales compared to the industry average. companies that manufacture electronic toys- t Like Games Inc. and Our Play Inc. Like Games was launched eight years ago, whereas y is a relatively new company that has been in operation for only the past two years. However, both companies have an equal market share with sales of $200,000 each. Youve collected company data to compare Like Games and Our Play. Last year, the average sales for all industry competitors was $510,000. As an anlyst, you want to make comments on the expected performance of these two companies in the coming year. Youve collected deta from the companies financial statements. This information is listed as follows Data Collected (in dollars) 5,400 110,000 90,000 Our Play ,800 160,000 industry Average 7,700 433,500 459,200 Net fixed assets Total assets Using this information, complete the folilowing stabements to include in your analysis. . Our Play has time to collect cash from its customers than it takes Like Games 2. Like Gamess fixed assets turnover ratic acquisition cost of its fxed assets Assuming that foxed assets prices (not book values) rose over the past six years due to days of sales ted up in receivables, which is much than the industry average. It takes Our Play than that of Our Play. This is because Like Games was formed eight years ago, so the s recorded at historic values when the company bought is assets and has been depreciated since then inflation, Our Play paid a e amount for its fixed 1. The average total assets turnover in the electronic toys industry is 1.09%, which means that $1.09 of sales is being generated with every dollar of investment in assets. A total assets turnover rati。ndates greater emciency. Both companies total assets turnover ratios than the industry average O Type here to search PrtScn
Answer #1
Quick ratio= (Cash and Equivalents + Marketable Securities + Accounts Receivable) / Current Liabilities
Quick ratio= (Current assets - Inventory) / Current liabilities
2= (56500-I)/19775
I= 16950
Inventory turnover ratio= COGS/ Average Inventory

Assuming, based on limited information, that the sales= COGS and average inventory levels remain stable

Inv TO= =200000/16950
11.7994 times
Industry average= 10.03 times

More the turnover, less the average inventory holding period and holding quantity. Hence option B is correct

Like games and Our play
1
Our Play Industry
days of sales in receivables= receivable/ sales * 365 receivable/ sales * 365
=7800/ 200000 * 365 =7700/ 510000 * 365
                               14.24          5.51
HIGHER than industry average
More time to collect cash
2
fixed asset TO ratio= net sales/ (Fixed assets - depreciation)
=200000/110000 =200000/160000
                                 1.82          1.25

Our Play was formed later when the asset prices had increased so it paid MORE than Like Games for acquiring FA

3
total asset TO ratio= net sales/ Total assets
=200000/190000 =200000/250000
                                 1.05          0.80

A GREATER total asset TO means greater efficiency since, more sales is generated for every $ of assets.
Both companies have a LOWER total asset TO than industry average